OpenStudy (mackenzie2013):

what is the theoretical probability over a coin flipped 200 times? ( in percentage )

OpenStudy (anonymous):

P = the probability of a success Q = 1 - P, the probability of failure n = the number of trials X = the number of successes Y = (X - nP)/sqrt(nPQ) and stick that in a normal distribution. using 104.5 as the number of successes, we get: (104.5-200*0.5/sqrt(200*0.5*0.5) = 0.6363... about 73.77 percent of the normal distribution is within that standard deviation, so your probability is about 26.22%